OpenStudy (anonymous):

If f(x)= (e^x)-(x^2)/2, then find the value of f(4)-f(0).

OpenStudy (campbell_st):

substitute x = 4 f(4) = (e^4 - 4^2)/2 f(0) = (e^0 - 0^2)/2 get the values then find the difference

OpenStudy (campbell_st):

or you could leave it in exact form

OpenStudy (anonymous):

I got that far, but i dont know what to do with the e^4

OpenStudy (anonymous):

If \[f(x)=e^x- \frac{ x^2 }{ 2 } \] then find the value of \[f(4)-f(0)\]

OpenStudy (freckles):

You can leave e^4 as e^4 or if you are allowed you can approximate it by using your calculator

OpenStudy (anonymous):

Oh. Okay thanks